OpenStudy (anonymous):

Can someone show me how to turn this parabola into standard form? y^2 + 4y - 8x + 4 = 0

OpenStudy (anonymous):

I'm not quite sure what the standard form of a parabola is

OpenStudy (whpalmer4):

standard form looks like \[y = ax^2+bx+c\]for a parabola that opens up or down and \[x = ay^2+by+c\]for a parabola that opens right or left

OpenStudy (anonymous):

So it would be 8x = y^2 + 4y + 4?

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is ?

OpenStudy (anonymous):

You asked me that in my last question and I think I do know what it is but you gave me a youtube video and I can't wait videos on this computer.

OpenStudy (anonymous):

*watch

OpenStudy (jdoe0001):

well, the video shows.... how to turn it into it =) thus

OpenStudy (anonymous):

My computer doesn't play videos though /:

OpenStudy (jdoe0001):

?

OpenStudy (whpalmer4):

\[8x=y^2+4y+4\]\[x=\frac{1}{8}(y^2+4y+4)\]Is your standard form. Now can you write the stuff in () as a square?

OpenStudy (anonymous):

The stuff in parenthesis as a square would be (y + 2)^2, right?

OpenStudy (anonymous):

so x = 1/8(y+2)^2 ?

OpenStudy (jdoe0001):

\(\bf y^2 + 4y - 8x + 4 = 0\implies 8x=(y^2 + 4y + 2^2)\\ \quad \\ \implies 8x=(y+2)^2\implies 8(x-0)=(y+2)^2\)

OpenStudy (anonymous):

Where did the 0 come from?

OpenStudy (jdoe0001):

8x = 8x+0 or 8x-0 = 8x

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

It can't be left at x = 1/8 (y + 2)^2 ?

OpenStudy (jdoe0001):

sure you can, then it won't be standard form per se

OpenStudy (anonymous):

I just went through my notes and my equation for horizontal parabolas is x = 1/4p y^2

OpenStudy (anonymous):

So I think that's the most my teacher expects from me for now lol

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

hmmm ahemm well.. you see... for a parabola either form will do \(\bf (y-k)^2=4p(x-h)\qquad x=ay^2+by+c\) are both valid standard forms of a parabola if you want the latter, just solve for "x" like whpalmer4 showed above if you want the former, then use that :)

OpenStudy (whpalmer4):

rearranging it as a square as you did is also the first step to putting it in "vertex" form: \[y = (x-h)^2+k\]where you can just read off the coordinates of the vertex as \((h,k)\)